Diffusion Indices. Suppose that there is a survey conducted among the purchasing managers to form a confidence index. The sample is composed of 150 respondents. Question is “What is your expectation about your sales over the next 3 months?”. Respondents can choose between three possible answers: “A. Increase”, “B. Decrease” and “C. Stay at the same level”. Responses are as follows: A: 37, B:43 and C:70. Normalized level for the confidence index is 100. Calculate the index value given the composition of responses.
Answer #1

Answer:

Index value of A

= Response of A * 100 / Total response

= 3700 / 150

= 24.67

Index value of B

= Response of B * 100 / Total response

= 4300 / 150

= 28.67

Index value of C

= Response of B * 100 / Total response

= 7000 / 150

= 46.67

Total of three index is 100